Solve by factoring.
x^2 + 2x – 48 = 0

Answer by funmath(2933)

Solve by factoring.
x^2 + 2x – 48 = 0
(x___)(x___)=0
Fill the blanks with two integers that multiply together to give you -48, but add together to give you +2:
(x+8)(x-6)=0
Set each factor = to 0 and solve for x:
x+8=0 or x-6=0
x+8-8=0-8 or x-6+6=0+6
x=-8 or x=6
Happy Calculating!!!